A recent quality control test at a local chocolate bar factory finds that the mass of the chocolate bars appear to be normally distributed with a mean mass of 52g and a standard deviation of 1.02g. What percentage of chocolate bars are between 49.96g and 53.02g?
Answer by Boreal(15235)

z=(x-mean)/sd
=(49.96-52)/1.02=-2
z=(53.02-52)/1.02=+1
probability of z between -2 and +1=0.8186 or 81.86%